# Long-time soliton dynamics via a coarse-grained space-time method

### v2

@@ VI.1 Massless case ($\kappa=0$)
Figure 9: Backreaction dynamics of the massless fermionic fluid in response to a capacitor with fixed charges ±$\pm Q$ at $\mp L/2$, with $Q=4$, $L=40$, and $g=1.2$. The dynamics is captured up to $T_{m}=25000$. The $x-t$ heatmaps at early times $0\leq t\leq 500$ for $\phi_{x}(x,t)$ and $\phi_{t}(x,t)$ are shown in (a) and (b), respectively. (c) shows the evolution of the current $J(t)$ at $x=100$ over the entire simulated time $0\leq t\leq T_{m}$, while (d) shows the current over an intermediate time $(t<1200)$, and (e) zooms into the current at early times $(t<120)$. The current at the center of the capacitor is shown in (f), while the electric field $E(t)$ there is plotted in (g). The field $\phi_{x}$ at $x=100$ is plotted in (h).
Through the bosonization of the Dirac field, the Schwinger process for massless fermions is described by the Klein-Gordon equation [34]
@@
which is simply Eq. 5 with $\kappa$ = 0. Recall that $\varphi$ is the real-valued scalar field and $F$ is the classical background field. The term proportional to $g$ engenders a mass gap in the bosonic sector that reflects the fact that the low-energy excitations of the fermionic theory are mesons, bound particle-antiparticle pairs of mass $e/\sqrt{\pi}$. This term was the first manifestation of quark trapping in a low-dimensional gauge theory, requiring a non-perturbative approach [27]. In modern parlance, the low-energy excitations of the massless Schwinger model are hybridized longitudinal excitations of the gauge field and fermionic dipoles.
As discussed in Eqs. 6, the current and charge densities of the fermionic fluid relates to the scalar field by $\rho=-g\partial_{x}\varphi$ and $J=g\partial_{t}\varphi$. Therefore, the edge fields $\phi_{x}$ and $\phi_{t}$, as defined in Eqs. 10 and 11 take on physical meanings as the average charge and average current integrated over a finite resolution $\Delta x$ and $\Delta t$, respectively. Hence, a discrete version of Eq. 24 written in terms of the fields $\phi_{x}$ and $\phi_{t}$ can be used to study the backreaction dynamics of the electron-positron fluid in the presence of the source. Before exploring the dynamics of massive fields, we benchmark our numerical method using this well-studied linear model of the Schwinger process [34].
The setup under consideration is the backreaction dynamics of massless fermions to a pair of fixed charges ±$\pm Q$ stationed at $\mp L/2$ that form a 1D capacitor  [^3] This setup has originally been studied using an axillary $\theta$ field to implement a non-zero external electrical field [27, 28]. We follow the source-field procedure used in Ref. [34] which is more general and flexible.. This setup creates an external electric field
@@
that sources the Klein-Gordon equation given in Eq. 24. The initial conditions are chosen to be $\varphi(x,0)=\partial_{t}\varphi(x,0)=0$ so that at $t=0$ there are no charges. We employ radiative boundary conditions - the details of which are discussed in Appendix C - which allow us to simulate long-time dynamics while simultaneously limiting the computational domain to a finite spatial region. It is also interesting to note that while we have no reference to topological BCs [34] the total charge within the system is still conserved at all times. This is due to the fact that the total initial charge is $Q=0$ and that DEC enforces exactly local charge conservation (to be discussed in details in Sec. VI.1.2). The DEC-QED simulations are shown in Fig. 9. The distance between the fixed charges is $L=40$, while $Q=4$ and $g=1.2$. In Fig. 9a, where the heatmap for $\phi_{x}(x,t)$ is plotted, we see that pairs of fermions are immediately created near the capacitor plates that screen the fixed charges ±$\pm Q$. Each capacitor is a source from which matter waves originate. These waves then travel within the light cone and create particle-antiparticle pairs on their path.
@@
The ability of DEC-QED to capture evolution over very long times while simultaneously probing the detailed dynamics that happen within much shorter time scales is critical to to the phenomenology of this problem. Figs. 9c, 9d, and 9e provide a snapshot of the current density at $x\!=\!100$ a finite distance to the right of the capacitor - in long, intermediate, and early times, respectively. From Fig. 9e, we can see that it takes a finite amount of time for the waves to reach this location. The first wave to arrive is the one that originates from the source at $+L/2$. Note that when it first appears at $x\!=\!100$, the oscillation has very high frequency which slows down while the amplitude increases. After some time, the wave sourced by the other charge at $-L/2$ arrives and interferes with the first wave. As just mentioned, because the frequencies of the waves also vary when they travel, the mismatch in frequencies of the two waves at $x\!=\!100$ results in a beating pattern seen at Fig. 9d. The widths of the beats stretch because both frequencies are decreasing over time. As seen in Fig. 9c, eventually at very long times both frequencies approach the asymptotic value and the beating pattern vanishes. The asymptotic carrier frequency of both waves, and hence of the overall oscillation frequency scale at every location, is precisely the effective mass $g$ (Fig. 10). Note that the rate at which the frequency approaches this asymptotic value sets the intermediate time scale of the backreaction dynamics, i.e. the time scale over which the beating pattern occurs.
@@
The envelope switches from beats to decay over longer time scales. This is expected because at asymptotically long times, it is known from other considerations [27] that the fixed charges will be totally screened and the leakage of current out of the capacitor will have to cease. The charge density also decays, as the plot for $\phi_{x}(t)$ in Fig. 9c indicates. The envelope decays at an asymptotic long-time rate of $1/\sqrt{t}$ [34], though this behavior emerges only well after the oscillations have settled into their asymptotic frequency.
@@
Within the capacitor region ($-L/2<x<L/2$), electron-positron pairs are rapidly created to screen the electric field generated by the fixed charges. Since the fermions are massless in this model, pair production occurs almost instantaneously, leading to fast, underdamped oscillations in the polarization current. Over time, these charge oscillations dissipate as the system radiates energy and approaches equilibrium. Both the current (Fig.,9f) and the electric field $E$ (Fig.,9g) gradually decay as the generated pairs effectively screen the sources.
@@
Figure 10: Oscillation frequency of the current $J$ at $x=100$ in the capacitor setup is plotted as a function of the period number. The parameters in the simulations are ±$\pm Q$ at $\mp L/2$, with $Q=4$, $L=40$, and $g=1.2$.
@@
Figure 11: Energy and charge conservation in the backreaction dynamics of the massless fluid in response to a capacitor with fixed charges. (a) The energy computed from the simulation results obtained from DEC is compared to those calculated using the Euler method and the Crank-Nicolson method. For all three methods, $\Delta x\!=\!0.333,\ \Delta t\!=\!0.125,T_{m}\!=\!1000$. (b) A closed-up view of the energy computed using the Euler method and using DEC. (c) The heat map of $\phi(f)$, evaluated at each primal face $f$ by summing up the edge fields $\phi_{x}$ and $\phi_{t}$ on the boundary of that face.
@@ VI.1.1 Conservation of energy
We also computed the energies associated with the backreaction dynamics and used it as a metric to compare DEC-QED with other approaches for simulating time dynamics. Fig. 11(a) shows the energy components as a function of time obtained from simulations using the Crank-Nicolson method, the Euler method, and DEC. Due to the underlying property of the Crank-Nicolson (CN) method in the way it takes time-averages of spatial derivatives, there is no consistent way to compute the energies from this method. As a result, the total energy is not conserved in the Crank-Nicolson method and decays over time (see Fig. 11a). Instead of the traditional time propagation, this method also requires matrix inversions to obtain the field at all locations at once, making it grossly inefficient and memory intensive when applied to large problems. To quote some numbers – to complete a simulation for $T_{m}=1000,\ L=40$ with grid spacings of $dt=0.125$ and $dx=0.333$, the CN method needs about one hour. The Euler method and DEC, on the other hand, both finish the simulation with the same parameters in under 5 seconds. These two methods also ensure that energy conservation is respected. The Euler method, however, introduces unphysical micro-oscillations in the total energy, as seen in Fig. 11b. These oscillations originate from the evaluation of the spatial and temporal derivatives using the discrete values of $\varphi$ obtained from solving the second-order Eq. 24. The errors are then carried over to the calculation of the Hamiltonian
[…]
@@ VI.1.2 Local conservation of charge
In the spacetime plane, the left-hand side of Eq. 28 is equivalent to ×$\nabla\!\times\!\nabla\varphi$. The discrete version of local charge conservation rule is obtained by integrating Eq. 28 over a primal face $f$ with dimensions $\Delta x$-by-$\Delta t$ and applying Stokes’ theorem
@@
Figure 13: Interaction between a fixed classical charge and an oppositely charged soliton in the context of the massive Schwinger process. The fixed charge is located at $x=0$. The initial condition for the scalar field $\varphi$ is the SG fluxon solution that is also initially placed at the origin with an initial velocity $u=0.55$. (a) and (b) show the resulting fields $\phi_{x}$ and $\phi_{t}$ for $g=0.1$ at early times ($2000<t<$ 200). The vertical green dashed line in each plot indicates the fixed location of the classical source. (c) and (d) also show the fields $\phi_{x}$ and $\phi_{t}$ at early times, but for $g=1.0$. (e) plots the field $\phi_{x}$ for $g=0.3$ at early times, while (f) plots the same field but at late times $5800\leq t\leq 6000$. (g) plots each of the terms in the Hamiltonian of the system and the total energy as functions of time. (h) plots the dependence on time of the energies within a finite region $-16\leq x\leq 16$. (i) shows the frequencies in the oscillations of each term in the Hamiltonian as a function of the period number.
@@ VI.2.1 Schwinger atom
We consider a setup composed of a single classical charge $Q$ fixed at the origin $x\!=\!0$. This source has to be balanced by an oppositely charged soliton in the scalar field so that the total electric energy is finite. The choice of initial conditions is therefore a delicate matter and we choose it to be the known analytical form for the kink solution of the unperturbed SG model (Eq. 19) that has a total charge of $2\pi g$. For the system to be charge-neutral, we choose $Q=-2\pi g$. At the boundaries of the finite computational domain, radiative BCs are applied to allow the mesons that are created during transient dynamics to leave the system without being artificially reflected back into the domain and affecting the true dynamics.
We first study the case where the initial position of the SG kink is $x_{0}\!=\!0$, exactly on top of the classical source, and has a starting velocity $u\!>\!0$. Due to the finite electric field generated in the region between the classical source and the soliton (Fig. 12a), the soliton is constantly pulled towards the source. As a result, a soliton with an initial velocity will oscillate around the source (Fig. 12b) and together they form a 1D “atom” which we refer to as the Schwinger atom. If we turn off this mediation of attractive interaction by removing the term $g^{2}\varphi$ from Eq. 30, we return to the pure SG equation on the left-hand side with a step function source $-gQ\Theta(x)$ on the right-hand side. The resulting dynamics of a soliton governed by this modified equation is shown in Fig. 12d, where instead of oscillating around the source, the charged soliton escapes towards $x\!\rightarrow\!-\infty$ at a constant velocity. It is helpful here to establish an analogy to fluxon dynamics in a JTL, which for the conditions stated here, has half of its length subject to a bias current (Fig. 12c). In that scenario, the dynamics in Fig. 12d is equivalent to that of a vortex initially at $x=0$ and moving into the biased region. In this region, the vortex experiences the Lorentz force from the bias current and is pulled towards negative $x$, causing it to slow down and switch its direction. Eventually the vortex exits the biased region and travels towards $x\!\rightarrow\!-\infty$.
[…]
@@
[…]
@@
Figure 15: Comparisons between DEC-QED and the Euler method in simulating the out-of-equilibrium dynamics of the massive Schwinger process when a soliton interacts with a fixed classical source. The classical charge is placed at $x\!=\!0$, while the soliton is initially at $x_{0}\!=\!-30$, away from the source and is moving towards it with velocity $u\!=\!0.55$. The 1+1D distribution of $\phi_{t}$ for $t\leq T_{m}$ is computed with $dx\!=\!0.02,\ dt\!=\!0.016$ using DEC and Euler’s method are shown in (a) and (b) respectively. The same field computed with $dx\!=\!0.1,\ dt\!=\!0.08$ using the two methods are shown in (c) and (d), respectively. $\phi_{t}(x)$ at the time slice $t=T_{m}$ obtained from both methods with $dx\!=\!0.02,\ dt\!=\!0.016$ are plotted in (e), while the results computed with $dx\!=\!0.1,\ dt\!=\!0.08$ are shown in (f) for DEC and in (g) for Euler.
@@
We show the evolution of $\phi_{t}$ over $3000\!\leq\!t\leq\!T_{m}\!=\!$ 300 computed with discretization steps $dx\!=\!0.02,\ dt\!=\!0.016$ using DEC-QED (Fig. 15a) and Euler (Figs. 15b). We see that with such a finely discretized grid, the two methods produce practically the same dynamics within the simulated time $t\!\leq\!T_{m}$. This is also evident in Fig.15e, where the details of $\phi_{t}$ at the last time slice $t\!=\!T_{m}$ obtained using the two methods are compared in the same plot. However, when the calculations are done with coarser discretization steps, we see that while DEC remains fairly accurate (Figs.15c, 15f), Euler’s method quickly diverges from the correct dynamics (Figs.15d, 15g).

### v3

@@ VI.1 Massless case ($\kappa=0$)
Figure 9: Backreaction dynamics of the massless fermionic fluid in response to a capacitor with fixed charges $\pm Q$ at $\mp L/2$, with $Q=4$, $L=40$, and $g=1.2$. The dynamics is captured up to $T_{m}=25000$. The xtx-t heatmaps at early times $0\leq t\leq 500$ for $\phi_{x}(x,t)$ and $\phi_{t}(x,t)$ are shown in (a) and (b), respectively. (c) shows the evolution of the current $J(t)$ at $x=100$ over the entire simulated time $0\leq t\leq T_{m}$, while (d) shows the current over an intermediate time $(t<1200)$, and (e) zooms into the current at early times $(t<120)$. The current at the center of the capacitor is shown in (f), while the electric field $E(t)$ there is plotted in (g). The field $\phi_{x}$ at $x=100$ is plotted in (h).
Through the bosonization of the Dirac field, the Schwinger process for massless fermions is described by the Klein-Gordon equation [chu2010capacitordischarge]
@@
which is simply Eq. 6 with $\kappa$ = 0. Recall that $\varphi$ is the real-valued scalar field and $F$ is the classical background field. The term proportional to $g$ engenders a mass gap in the bosonic sector that reflects the fact that the low-energy excitations of the fermionic theory are mesons, bound particle-antiparticle pairs of mass $e/\sqrt{\pi}$. This term was the first manifestation of quark trapping in a low-dimensional gauge theory, requiring a non-perturbative approach [coleman1975]. In modern parlance, the low-energy excitations of the massless Schwinger model are hybridized longitudinal excitations of the gauge field and fermionic dipoles.
As discussed in Eqs. 7, the current and charge densities of the fermionic fluid relates to the scalar field by $\rho=-g\partial_{x}\varphi$ and $J=g\partial_{t}\varphi$. Therefore, the edge fields $\phi_{x}$ and $\phi_{t}$, as defined in Eqs. 11 and 12 take on physical meanings as the average charge and average current integrated over a finite resolution $\Delta x$ and $\Delta t$, respectively. Hence, a discrete version of Eq. 25 written in terms of the fields $\phi_{x}$ and $\phi_{t}$ can be used to study the backreaction dynamics of the electron-positron fluid in the presence of the source. Before exploring the dynamics of massive fields, we benchmark our numerical method using this well-studied linear model of the Schwinger process [chu2010capacitordischarge].
The setup under consideration is the backreaction dynamics of massless fermions to a pair of fixed charges $\pm Q$ stationed at $\mp L/2$ that form a 1D capacitor  [^3] This setup has originally been studied using an axillary $\theta$ field to implement a non-zero external electrical field [coleman1975, coleman1976]. We follow the source-field procedure used in Ref. [chu2010capacitordischarge] which is more general and flexible.. This setup creates an external electric field
@@
that sources the Klein-Gordon equation given in Eq. 25, where $\Theta(x)$ represents a Heaviside step function. The initial conditions are chosen to be $\varphi(x,0)=\partial_{t}\varphi(x,0)=0$ so that at $t=0$ there are no charges. We employ radiative boundary conditions - the details of which are discussed in Appendix C - which allow us to simulate long-time dynamics while simultaneously limiting the computational domain to a finite spatial region. It is also interesting to note that while we have no reference to topological BCs [chu2010capacitordischarge] the total charge within the system is still conserved at all times. This is due to the fact that the total initial charge is $Q=0$ and that DEC enforces exactly local charge conservation (to be discussed in details in Sec. VI.1.2). The DEC-QED simulations are shown in Fig. 9. The distance between the fixed charges is $L=40$, while $Q=4$ and $g=1.2$. In Fig. 9a, where the heatmap for $\phi_{x}(x,t)$ is plotted, we see that pairs of fermions are immediately created near the capacitor plates that screen the fixed charges $\pm Q$. Each capacitor is a source from which matter waves originate. These waves then travel within the light cone and create particle-antiparticle pairs on their path.
@@
The ability of DEC-QED to capture evolution over very long times while simultaneously probing the detailed dynamics that happen within much shorter time scales is critical to to the phenomenology of this problem. Figs. 9c, 9d, and 9e provide a snapshot of the current density at x=100x\!=\!100 - a finite distance to the right of the capacitor - in long, intermediate, and early times, respectively. From Fig. 9e, we can see that it takes a finite amount of time for the waves to reach this location. The first wave to arrive is the one that originates from the source at +L/2+L/2. Note that when it first appears at x=100x\!=\!100, the oscillation has very high frequency which slows down while the amplitude increases. After some time, the wave sourced by the other charge at L/2-L/2 arrives and interferes with the first wave. As just mentioned, because the frequencies of the waves also vary when they travel, the mismatch in frequencies of the two waves at x=100x\!=\!100 results in a beating pattern seen at Fig. 9d. The widths of the beats stretch because both frequencies are decreasing over time. As seen in Fig. 9c, eventually at very long times both frequencies approach the asymptotic value and the beating pattern vanishes. The asymptotic carrier frequency of both waves, and hence of the overall oscillation frequency scale at every location, is precisely the effective mass $g$ (Fig. 10). Note that the rate at which the frequency approaches this asymptotic value sets the intermediate time scale of the backreaction dynamics, i.e. the time scale over which the beating pattern occurs.
@@
The envelope switches from beats to decay over longer time scales. This is expected because at asymptotically long times, it is known from other considerations [coleman1975] that the fixed charges will be totally screened and the leakage of current out of the capacitor will have to cease. The charge density also decays, as the plot for $\phi_{x}(t)$ in Fig. 9c indicates. The envelope decays at an asymptotic long-time rate of $1/\sqrt{t}$ [chu2010capacitordischarge], though this behavior emerges only well after the oscillations have settled into their asymptotic frequency.
@@
Within the capacitor region (L/2<x<L/2-L/2<x<L/2), electron-positron pairs are rapidly created to screen the electric field generated by the fixed charges. Since the fermions are massless in this model, pair production occurs almost instantaneously, leading to fast, underdamped oscillations in the polarization current. Over time, these charge oscillations dissipate as the system radiates energy and approaches equilibrium. Both the current (Fig.,9f) and the electric field $E$ (Fig.,9g) gradually decay as the generated pairs effectively screen the sources.
@@
Figure 10: Oscillation frequency of the current JJ at $x=100$ in the capacitor setup is plotted as a function of the period number. The parameters in the simulations are $\pm Q$ at $\mp L/2$, with $Q=4$, $L=40$, and $g=1.2$.
Eneg
@@
Figure 11: Energy and charge conservation in the backreaction dynamics of the massless fluid in response to a capacitor with fixed charges. (a) The energy computed from the simulation results obtained from DEC is compared to those calculated using the Euler method and the Crank-Nicolson method. For all three methods, $\Delta x\!=\!0.333,\ \Delta t\!=\!0.125,T_{m}\!=\!1000$. (b) A closed-up view of the energy computed using the Euler method and using DEC. (c) The heat map of $\varphi(f)$, evaluated at each primal face $f$ by summing up the edge fields $\phi_{x}$ and $\phi_{t}$ on the boundary of that face.
@@ VI.1.1 Conservation of energy
We also computed the energies associated with the backreaction dynamics and used it as a metric to compare DEC-QED with other approaches for simulating time dynamics. Fig. 11(a) shows the energy components as a function of time obtained from simulations using the Crank-Nicolson method, the Euler method, and DEC. Due to the underlying property of the Crank-Nicolson (CN) method in the way it takes time-averages of spatial derivatives, there is no consistent way to compute the energies from this method. As a result, the total energy is not conserved in the Crank-Nicolson method and decays over time (see Fig. 11a). Instead of the traditional time propagation, this method also requires matrix inversions to obtain the field at all locations at once, making it less efficient and more memory intensive when applied to large problems. One should note, however, that the CN method can be further optimized to achieve near $O(N)-O(N\text{log}N)$ scaling. For very long-time simulations, though, the runtime differences between CN, Euler’s method, and our approach eventually become significant. Euler’s method and our spacetime-DEC approach also ensure that energy conservation is respected. Euler’s method, however, introduces unphysical micro-oscillations in the total energy, as seen in Fig. 11b. These oscillations originate from the evaluation of the spatial and temporal derivatives using the discrete values of $\varphi$ obtained from solving the second-order Eq. 25. The errors are then carried over to the calculation of the Hamiltonian
$H(t)=\int_{-L/2}^{L/2}dx\frac{1}{2}\left[\left(\partial_{x}\varphi\right)^{2}+\left(\partial_{t}\varphi\right)^{2}+\left(g\varphi+F\right)^{2}\right],$ (27)
@@ VI.1.2 Local conservation of charge
In the spacetime plane, the left-hand side of Eq. 29 is equivalent to $\nabla\!\times\!\nabla\varphi$. The discrete version of local charge conservation rule is obtained by integrating Eq. 29 over a primal face $f$ with dimensions $\Delta x$-by-$\Delta t$ and applying Stokes’ theorem
@@
Figure 13: Interaction between a fixed classical charge and an oppositely charged soliton in the context of the massive Schwinger process. The fixed charge is located at $x=0$. The initial condition for the scalar field $\varphi$ is the SG fluxon solution that is also initially placed at the origin with an initial velocity $u=0.55$. (a) and (b) show the resulting fields $\phi_{x}$ and $\phi_{t}$ for $g=0.1$ at early times ($0<t<200$). The vertical green dashed line in each plot indicates the fixed location of the classical source. (c) and (d) also show the fields $\phi_{x}$ and $\phi_{t}$ at early times, but for $g=1.0$. (e) plots the field $\phi_{x}$ for $g=0.3$ at early times, while (f) plots the same field but at late times $5800\leq t\leq 6000$. (g) plots each of the terms in the Hamiltonian of the system and the total energy as functions of time. (h) plots the dependence on time of the energies within a finite region $16\leq x\leq 16$. (i) shows the frequencies in the oscillations of each term in the Hamiltonian as a function of the period number.
@@ VI.2.1 Schwinger atom
We consider a setup composed of a single classical charge $Q$ fixed at the origin x=0x\!=\!0. This source has to be balanced by an oppositely charged soliton in the scalar field so that the total electric energy is finite. The choice of initial conditions is therefore a delicate matter and we choose it to be the known analytical form for the kink solution of the unperturbed SG model (Eq. 20) that has a total charge of $2\pi g$. For the system to be charge-neutral, we choose $Q=-2\pi g$. At the boundaries of the finite computational domain, radiative BCs are applied to allow the mesons that are created during transient dynamics to leave the system without being artificially reflected back into the domain and affecting the true dynamics.
We first study the case where the initial position of the SG kink is $x_{0}\!=\!0$, exactly on top of the classical source, and has a starting velocity u>0u\!>\!0. Due to the finite electric field generated in the region between the classical source and the soliton (Fig. 12a), the soliton is constantly pulled towards the source. As a result, a soliton with an initial velocity will oscillate around the source (Fig. 12b) and together they form a 1D “atom” which we refer to as the Schwinger atom. If we turn off this mediation of attractive interaction by removing the term $g^{2}\varphi$ from Eq. 31, we return to the pure SG equation on the left-hand side with a step function source $gQ\Theta(x)$ on the right-hand side. The resulting dynamics of a soliton governed by this modified equation is shown in Fig. 12d, where instead of oscillating around the source, the charged soliton escapes towards $x\!\rightarrow\!$-$\infty$ at a constant velocity. It is helpful here to establish an analogy to fluxon dynamics in a JTL, which for the conditions stated here, has half of its length subject to a bias current (Fig. 12c). In that scenario, the dynamics in Fig. 12d is equivalent to that of a vortex initially at $x=0$ and moving into the biased region. In this region, the vortex experiences the Lorentz force from the bias current and is pulled towards negative $x$, causing it to slow down and switch its direction. Eventually the vortex exits the biased region and travels towards $x\!\rightarrow\!$-$\infty$.
Figure 14: Interaction between a fixed classical charge and an oppositely charged soliton. The classical charge is placed at $x=0$, while the soliton is initially at a distance away from the source and is moving towards it. (a), (b), (c), and (d) show the dynamics of the charge current $\phi_{t}(x,t)$ at early times $0\leq t\leq 200$ for g=0.1, 0.3, 0.35,g=0.1,\ 0.3,\ 0.35, and $0.4$, respectively. In these figures the initial location of the soliton is $x_{0}=-30$, and vertical green dashed line indicates the fixed location of the classical source. (e) plots the early-time dynamics of $\phi_{x}$ for $g=0.3$ and $x_{0}=-10$, and (f) plots the late-time dynamics of the same field. (g) shows the frequencies in the oscillations of each term in the Hamiltonian as a function of the period number. (h) plots the dependence on time of the energies within a finite region $16\leq x\leq 16$.
@@
We next consider a situation where the soliton is initially at a distance away from the source and is moving towards it. We again choose the initial condition to be a SG soliton with initial position $x_{0}$ and velocity $u=0.55$. In Figs. 14a-d the results for the current $\phi_{t}(x,t)$ at early times $0\leq t\leq 200$ are shown for g=0.1, 0.3, 0.35,g=0.1,\ 0.3,\ 0.35, and $0.4$, respectively. In these figures, the soliton is initially located at $x_{0}=-30$. As the coupling strength $g$ increases, the shape and trajectory of the SG soliton become increasingly distorted and it emits radiation more rapidly. The soliton accelerates as it approaches the external charge, as reflected in the changing slope of its trajectory — most clearly seen in Fig. 14a. However, its speed remains bounded by the speed of light, and in Fig. 14b, we observe the soliton approaching this relativistic limit. In Figs. 14b and 14c, we also see that as the soliton reaches its maximum distance from the source and begins to reverse direction due to the attractive interaction, part of its charge separates and continues to propagate away. Throughout its motion, the soliton gradually loses its integrity, continuously shedding radiation in the form of mesons that travel outward from the source-soliton system.
@@
Figure 15: Comparisons between DEC-QED and the Euler method in simulating the out-of-equilibrium dynamics of the massive Schwinger process when a soliton interacts with a fixed classical source. The classical charge is placed at x=0x\!=\!0, while the soliton is initially at $x_{0}\!=\!-30$, away from the source and is moving towards it with velocity u=0.55u\!=\!0.55. The 1+1D distribution of $\phi_{t}$ for $t\leq T_{m}$ is computed with dx=0.02,dt=0.016dx\!=\!0.02,\ dt\!=\!0.016 using DEC and Euler’s method are shown in (a) and (b) respectively. The same field computed with dx=0.1,dt=0.08dx\!=\!0.1,\ dt\!=\!0.08 using the two methods are shown in (c) and (d), respectively. $\phi_{t}(x)$ at the time slice $t=T_{m}$ obtained from both methods with dx=0.02,dt=0.016dx\!=\!0.02,\ dt\!=\!0.016 are plotted in (e), while the results computed with dx=0.1,dt=0.08dx\!=\!0.1,\ dt\!=\!0.08 are shown in (f) for DEC and in (g) for Euler.
@@
We show the evolution of $\phi_{t}$ over $0\!\leq\!t\leq\!T_{m}\!=\!300$ computed with discretization steps dx=0.02,dt=0.016dx\!=\!0.02,\ dt\!=\!0.016 using DEC-QED (Fig. 15a) and Euler (Figs. 15b). We see that with such a finely discretized grid, the two methods produce practically the same dynamics within the simulated time $t\!\leq\!T_{m}$. This is also evident in Fig.15e, where the details of $\phi_{t}$ at the last time slice $t\!=\!T_{m}$ obtained using the two methods are compared in the same plot. However, when the calculations are done with coarser discretization steps, we see that while DEC remains fairly accurate (Figs.15c, 15f), Euler’s method quickly diverges from the correct dynamics (Figs.15d, 15g).